OpenStudy (anonymous):

A 16 foot ladder leans aginst a wall at 68 degrees angle of elevation. Find the distance from the ladders base to the wall?

OpenStudy (cwrw238):

cos 68 = d / 16 solve for d

OpenStudy (anonymous):

4.25 rounded would be 4.6?

OpenStudy (cwrw238):

4.3

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

hold on wrong answer d = 16 * cos 68 = 16*0.3746 = 6.0